# Optimal estimation for regression discontinuity design

### v1

@@ Abstract
We develop a finite-sample optimal estimator for regression discontinuity designs when the outcomes are bounded, including binary outcomes as the leading case. Our finite-sample optimal estimator achieves the exact minimax mean squared error among linear shrinkage estimators with nonnegative weights when the regression function of a bounded outcome lies in a Lipschitz class. Although the original minimax problem involves an iterating (nn+1)-dimensional non-convex optimization problem where $n$ is the sample size, we show that our estimator is obtained by solving a convex optimization problem. A key advantage of our estimator is that the Lipschitz constant is the only tuning parameter. We also propose a uniformly valid inference procedure without a large-sample approximation. In a simulation exercise for small samples, our estimator exhibits smaller mean squared errors and shorter confidence intervals than conventional large-sample techniques which may be unreliable when the effective sample size is small. We apply our method to an empirical multi-cutoff design where the sample size for each cutoff is small. In the application, our method yields informative confidence intervals, in contrast to the leading large-sample approach.
@@
T1 The study was supported by JSPS KAKENHI Grant Numbers JP22K13373 (Ishihara), and JP21K13269 (Sawada). We thank Yu-Chang Chen, Atsushi Inoue, Timothy Neal, Michal Kolesár, Soonwoo Kwon and Ke-Li Xu, as well as seminar participants at Japanese Joint Statistical Meeting, Hitotsubashi University, Kansai Keiryo Keizaigaku Kenkyukai and Tohoku-NTU Joint Seminar, Econometric Society World Congress 2025 for insightful comments. T2First Version: September 23, 2025
@@ 1 Introduction
Large-sample approximation is the basis for the leading estimators for regression discontinuity (RD) designs (Imbens and Kalyanaraman, 2012; Calonico et al., 2014, for example). RD designs involve the estimation of conditional expectation functions at a cutoff point on the support of a running variable. Hence, the effective observations are limited to the neighborhood of the cutoff, and the number of these observations can be small even if the total sample size is large (Cattaneo et al., 2015; Canay and Kamat, 2017). For example, the effective sample can be small for designs with multiple cutoffs, with a cutoff at the tail of the distribution, or with subgroup analyses. In small samples, the large-sample asymptotics may not provide good approximations of the behaviors of the existing estimators, and hence, their stated desirable properties may be lost.
@@
A few studies consider finite-sample minimax estimators for RD designs. [^1] Throughout the manuscript, we compare our estimator with existing finite-sample minimax estimators. Another notable approach is a finite-sample valid estimation and inference based on the local randomization of the RD design (Cattaneo et al., 2015, 2016, 2017). The local randomization approach is based on an assumption that the running variable is randomly assigned with a constant regression function within a given small window around the threshold (Cattaneo et al., 2024b), while we consider a smooth but nonconstant regression function within the window. For example, Armstrong and Kolesár (2018) and Imbens and Wager (2019) propose finite-sample minimax linear estimators under smoothness of the regression function. However, these minimax estimators require the knowledge of the conditional variance function, which is unknown in practice. While the variance can be estimated, we cannot guarantee the theoretical validity of the plug-in estimators with the estimated variance in finite samples. Furthermore, the construction of finite-sample valid confidence intervals based on these estimators additionally requires the normality of the regression errors.
@@
In this study, we propose finite-sample estimation and inference methods for RD designs with binary outcomes. For a binary dependent variable, all features of its conditional distribution, including its conditional variance, are a known function of its conditional mean function. We establish the finite-sample validity of our methods under a smoothness restriction on the conditional mean function, taking into account the implicit restrictions it imposes on the entire conditional distribution. In other words, our procedure is both feasible and theoretically valid without either the knowledge or estimation of the conditional variance, or more generally, any features of the conditional distribution except the smoothness of the conditional mean.
@@
More specifically, we consider a minimax optimal estimator among a class of linear shrinkage estimators for the regression function at a boundary point where the regression function satisfies the Lipschitz continuity. The class of linear shrinkage estimators is of the form $\sum_{i=1}^{n}w_{i}(Y_{i}-1/2)+1/2$ with $\sum_{i=1}^{n}w_{i}\leq 1$ and $w_{i}\geq 0$, where $Y_{1},...,Y_{n}$ are the observed outcomes on either side of the boundary. The shrinkage toward $1/2$ is motivated by the fact that the regression function is bounded and takes values in $[0,1]$, leading to a scope of efficiency gain by shrinkage. Given the class of linear shrinkage estimators, we derive a linear shrinkage estimator that minimizes the maximum mean squared error (MSE) under the Lipschitz continuity with a known Lipschitz constant. In other words, we assume the researcher’s a priori knowledge of the bound on how much the function value can change if the running variable is changed by one unit. We emphasize that this Lipschitz constant is the only tuning parameter. Furthermore, we show that the minimax estimator is the solution to a convex optimization problem, which is computationally feasible. Hence, we provide a practical exact finite-sample estimator when the outcome is binary.
@@
Our estimator is widely applicable to many practical RD designs. Binary outcomes are one of the most common types in empirical applications. For example, the following outcome variables are all binary: an indicator for winning the next election in the famous U.S. House election study by Lee (2008); a corruption indicator in Brollo et al. (2013); a mortality indicator in Card et al. (2009); and indicators for student’s enrollment and dropout in Melguizo et al. (2016) and Cattaneo et al. (2021). Furthermore, the first stage in fuzzy RD designs often involves a treatment status as the binary dependent outcome. Moreover, the minimax optimality of our estimator for binary outcomes immediately extends to that for bounded outcomes because the variance of any linear estimator is maximized when the outcomes are Bernoulli given the conditional mean function. Hence, our estimator can be applied not only to the binary-outcome case but also to the bounded-outcome case. As a result, our estimator is a practical finite-sample estimation method for frequently used outcome variables in RD designs.
@@
Our method also complements existing minimax estimators. We compare our estimator to a version of the existing minimax estimators (Armstrong and Kolesár, 2018; Imbens and Wager, 2019) and demonstrate that our method has better finite-sample performance than the existing approach while their asymptotic behaviors are similar. Specifically, we consider a minimax linear estimator obtained under a misspecified model where the conditional mean and variance are unrelated, the variance is known, and the regression function lies in a Lipschitz class with no bounds on function values. This estimator is not directly feasible in our binary-outcome setting, in which the variance is unknown. As a feasible version of this estimator, we consider the one obtained under the assumption of constant variance of $1/4$, which is the maximum possible variance of a binary variable. For binary outcomes, we theoretically show that the efficiency gain from our estimator relative to the above alternative estimator tends to vanish as the sample size increases. Nevertheless, for small samples, we numerically demonstrate that the alternative method can result in a $5\%$ to $20\%$ increase in the worst-case root MSE due to model misspecification. Hence, our method supplements the existing minimax estimators with better finite-sample performance and similar asymptotic behaviors in a binary-outcome setting.
@@
We also propose confidence intervals that have correct coverage in finite samples uniformly over the Lipschitz class. We construct the confidence intervals by inverting one-sided or two-sided uniformly valid tests that use a linear estimator as a test statistic. To construct a uniformly valid test, we propose a simulation-based approximation to the distribution of the test statistic by drawing samples from a multivariate Bernoulli distribution satisfying the null restriction. We then numerically optimize the critical value so that the worst-case rejection probability is equal to or smaller than the significance level. A computational challenge with this approach is the calculation of the worst-case rejection probability, which involves an optimization over an $(n+1)$-dimensional parameter. We overcome this challenge by deriving a simple characterization of the worst-case rejection probability under the Lipschitz continuity, which significantly reduces the computational burden. We also emphasize that our confidence intervals are valid in finite samples for binary outcomes. This is in contrast to existing inference methods that are based on either a large-sample approximation or the restrictive assumption of Gaussian errors with a known variance.
@@
The same inference approach does not apply to bounded outcomes because the simple characterization of the worst-case rejection probability relies on the fact that the outcome is binary. For bounded outcomes, we provide an alternative finite-sample inference procedure based on a uniform bound on the rejection probability obtained by the Hoeffding’s inequality. The resulting confidence intervals have correct coverage in finite samples but can be conservative like usual Hoeffding’s-inequality-based confidence intervals in other contexts.
@@
We demonstrate the performance of our methods through simulations and an empirical application. In simulations, our estimator achieves substantially small MSEs relative to the leading large-sample estimators when the sample size is small. Furthermore, our estimator has a similar behavior to the large-sample estimators when the sample size is larger; the differences in the MSE shrink as the number of observations increases. Our proposed inference method also achieves guaranteed coverage rates with shorter confidence intervals when the sample size is small. Hence, our estimator is optimal in theory and useful in practice.
@@
We illustrate our methods by revisiting Brollo et al. (2013), who estimate the impact of additional government revenues on corruption. They exploit a regional fiscal rule in Brazil, where federal transfers to municipal governments change exogenously at given population thresholds. This setting is a multi-cutoff RD design with a small sample size near each cutoff. We demonstrate that our estimates are similar to the conventional estimates for the large sample pooling multiple cutoffs. Nevertheless, our inference method gives much shorter confidence intervals than the conventional methods when we focus on a small sample near each cutoff. As a result, our estimates provide more informative results than the estimates from the conventional methods.
@@
Both simulations and application results indicate that the finite-sample estimations are challenging while our estimator has a potential to provide informative estimates. Hence, our estimator is a practical last resort for an empirical researcher who faces a research question with a small effective sample size for RD designs.
@@
In addition to the contributions to estimation in RD designs, we contribute to the vast literature on minimax estimation. Donoho (1994) considers minimax affine estimation and inference on linear functionals in nonparametric regression models with Gaussian errors. Recently, his framework has been applied to estimation and inference on treatment effects in a variety of settings, including RD designs (Armstrong and Kolesár, 2018; Gao, 2018; Imbens and Wager, 2019; Kwon and Kwon, 2020; Armstrong and Kolesár, 2021; de Chaisemartin, 2021; Rambachan and Roth, 2023). We complement these existing studies by studying nonparametric regression models with Bernoulli dependent variables, which are not covered by their frameworks. To the best of our knowledge, no general minimax estimator under squared error loss is established for the problem of estimating linear functionals in this setting. [^2] DeRouen and Mitchell (1974) derives a $\Gamma$-minimax estimator for a linear combination of the success probabilities of multiple independent binomial variables when the class of prior distributions consists of distributions with the same, known means. No solution is known even for the estimation of the difference in the success probability between two independent binomial variables of unequal numbers of trials (Lehmann and Casella, 1998, Example 5.1.9). [^3] For the estimation of the success probability of a single binomial variable, a linear shrinkage (toward $1/2$) estimator is minimax among all estimators (Lehmann and Casella, 1998, Example 5.1.7). Marchand and MacGibbon (2000) consider this problem with a restricted parameter space. They show that, when the success probability is known to lie in a symmetric interval around $1/2$, a linear shrinkage estimator is minimax among all linear estimators. We contribute to this underexplored literature by developing a minimax estimator for a regression function at a point, a particular linear functional, within the class of linear shrinkage estimators under the Lipschitz continuity of the regression function.
@@ 2 Our minimax estimator and its properties
RD designs exploit a discontinuous change in the treatment status when a running variable exceeds a cutoff point. For example, Brollo et al. (2013) exploit discontinuous increases in the amount of central government subsidy for a local government when its residing population equals or exceeds a threshold level. The target parameter of a RD design is the average treatment effect at the cutoff point and it is identified as the difference in conditional expectation functions evaluated at the cutoff point. Hence, its estimation involves the nonparametric estimation of the conditional mean functions at their boundary point.
@@ 2.1 Setting
Suppose that we have a random sample $\{Y_{i},D_{i},R_{i}\}_{i=1}^{N}$, where $R_{i}\in\mathbb{R}^{d_{r}}$ is a $d_{r}(\geq 1)$-dimensional vector of running variables, $Y_{i}$ is a binary outcome, $D_{i}$ is a binary treatment assigned as $D_{i}=1\{R_{i}\in\mathcal{T}\}$, and $\mathcal{T}\subset\mathbb{R}^{d_{r}}$ is a known treated region. The leading case is the one where $R_{i}$ is univariate ($d_{r}=1$) and $\mathcal{T}=[c,\infty)$ for some known cutoff $c$, but the following arguments apply to a multidimensional case (i.e., $d_{r}>1$) as well. Suppose
@@
for some unknown function $f:\{0,1\}\times\mathbb{R}^{d_{r}}\rightarrow[0,1]$. Let $R_{0}$ be a fixed boundary point of the treatment region $\mathcal{T}$. When $f(d,r)$ represents the conditional expectation function of the underlying potential outcome $Y_{i}(d)$ conditional on $R_{i}=r$ for each $d\in\{0,1\}$, $f(1,R_{0})-f(0,R_{0})$ is interpreted as the average treatment effect at the boundary point $R_{0}$ (Hahn et al., 2001). The data $\{Y_{i},D_{i},R_{i}\}_{i=1}^{N}$ can be divided into $\{Y_{i,+},R_{i,+}\}_{i=1}^{n_{+}}$ and $\{Y_{i,-},R_{i,-}\}_{i=1}^{n_{-}}$, where the former is the data from the treatment group and the latter is the data from the control group. We use the two samples separately to estimate $f(1,R_{0})$ and $f(0,R_{0})$, respectively.
Without loss of generality, we consider the estimation of $f(1,R_{0})$ throughout this section, except in Remark 2.5 at the end of this section where we discuss the estimation of $f(1,R_{0})-f(0,R_{0})$. To simplify the notation, we use $\{Y_{i},R_{i}\}_{i=1}^{n}$ to denote $\{Y_{i,+},R_{i,+}\}_{i=1}^{n_{+}}$, so that $R_{i}\in{\cal T}$ for all i=1,ni=1,...n. Furthermore, we use $f(\cdot)$ to denote $f(1,\cdot)$. Additionally, our analysis conditions on the realization of $\{R_{i}\}_{i=1}^{n}$, and we treat $\{R_{i}\}_{i=1}^{n}$ as deterministic, so that $P(Y_{i}=1)=f(R_{i})$ for all $i=1,\ldots,n$. Let $p_{i}\equiv f(R_{i})$ for $i=0,1,\ldots,n$ and $\bm{p}\equiv(p_{0},p_{1},\ldots,p_{n})^{\prime}\in[0,1]^{n+1}$. Without loss of generality, we assume that $R_{0}=0$ and $\|R_{0}\|\leq\|R_{1}\|\leq\cdots\leq\|R_{n}\|$, where $\|\cdot\|$ is a norm on $\mathbb{R}^{d_{r}}$. The following theoretical result holds for any norm, but we focus on the Euclidean norm in numerical exercises, simulations, and the empirical application.
For the parameter of interest $p_{0}=f(0)$, we consider the following linear shrinkage estimator:
@@
where $\mathcal{W}\equiv\left\{\bm{w}\in\mathbb{R}^{n}:\sum_{i=1}^{n}w_{i}\leq 1\ \text{and}\ w_{i}\geq 0\ \text{for all $i$}\right\}$. When $\sum_{i=1}^{n}w_{i}=1$, $\hat{p}_{0}(\bm{w})=\sum_{i=1}^{n}w_{i}Y_{i}$, and there is no shrinkage. When $\sum_{i=1}^{n}w_{i}<1$, $\hat{p}_{0}(\bm{w})$ is an estimator that shrinks toward $1/2$.
We assume that $f$ lies in the Lipschitz class
@@
where $C$ denotes the Lipschitz constant. This assumption implies that $\bm{p}\in[0,1]^{n+1}$ satisfies $|p_{i}-p_{j}|\leq C\|R_{i}-R_{j}\|$ for all $i$ and $j$. Conversely, if $|p_{i}-p_{j}|\leq C\|R_{i}-R_{j}\|$ for all $i$ and $j$, we can find a function $f\in\mathcal{F}_{\text{Lip}}(C)$ such that $f(R_{i})=p_{i}$ for all $i$ (Beliakov, 2006). Hence, the parameter space of $\bm{p}$ can be written as follows:
@@
To simplify the expression in (2.3), we redefine $p_{i}$ as $\theta_{i}\equiv p_{i}-1/2$ for $i=0,1,\ldots,n$ and let $\bm{\theta}\equiv(\theta_{0},\theta_{1},\ldots,\theta_{n})^{\prime}$, so that the problem is
@@
Hence, we obtain the weight vector that minimizes the maximum MSE by solving (2.4).
@@
The class of linear shrinkage estimators (2.1) eliminates linear estimators with negative weights. Hence it excludes the local polynomial estimators, which are commonly employed in RD designs. Nevertheless, the linear minimax MSE estimator has nonnegative weights in related setups where an outcome is non-binary (e.g. Gaussian outcomes) and its regression function lies in the Lipschitz class with a known conditional variance: see Section 3 and Appendix D. Hence, we focus on linear shrinkage estimators with nonnegative weights.

### v2

@@ Abstract
We develop a finite-sample optimal estimator for regression discontinuity design when the outcomes are bounded, including binary outcomes as the leading case. Our estimator achieves minimax mean squared error among linear shrinkage estimators with nonnegative weights when the regression function lies in a Lipschitz class. Although the original minimax problem involves an iterative noncovex optimization problem, we show that our estimator is obtained by solving a convex optimization problem. A key advantage of the proposed estimator is that the Lipschitz constant is its only tuning parameter. We also propose a uniformly valid inference procedure without a large-sample approximation. In a simulation exercise for small samples, our estimator exhibits smaller mean squared errors and shorter confidence intervals than those of conventional large-sample techniques. In an empirical multi-cutoff design in which the sample size for each cutoff is small, our method yields informative confidence intervals, in contrast to the leading large-sample approach.
@@
T1 This study was supported by JSPS KAKENHI Grant Numbers JP22K13373 (Ishihara) and JP21K13269 (Sawada). We thank Yu-Chang Chen, Atsushi Inoue, Timothy Neal, Michal Kolesár, Soonwoo Kwon, Tomasz Olma, and Ke-Li Xu, as well as seminar participants at the Japanese Joint Statistical Meeting, Hitotsubashi University, Kansai Keiryo Keizaigaku Kenkyukai, the Tohoku-NTU Joint Seminar, the Econometric Society World Congress 2025, and LMU-Todai Econometrics Workshop, for their insightful comments. T2First Version: September 23, 2025; Current Version: December 26, 2025
@@ 1 Introduction
A large-sample approximation is the basis for the leading estimators for regression discontinuity (RD) designs (Imbens.Kalyanaraman2012; Calonico.Cattaneo.Titiunik2014, for example). RD designs involve the estimation of conditional expectation functions at a cutoff point on the support of a running variable. Hence, effective observations are limited to the neighborhood of the cutoff, and the number of these observations can be small, even if the total sample size is large (Cattaneo.Frandsen.Titiunik2015; Canay.Kamat2017). For example, an effective sample can be small for designs with multiple cutoffs, with a cutoff at the tail of the distribution, or with subgroup analyses. In small samples, the large-sample asymptotics may not provide good approximations of the behaviors of the existing estimators; hence, their desirable properties may be lost.
@@
Some studies have considered finite-sample minimax estimators for RD designs. [^1] Throughout the manuscript, we compare our estimator with existing finite-sample minimax estimators. Another notable approach is finite-sample valid estimation and inference based on the local randomization of the RD design (Cattaneo.Frandsen.Titiunik2015; cattaneoInferenceRegressionDiscontinuity2016; cattaneoComparingInferenceApproaches2017). The local randomization approach is based on the assumption that the running variable is randomly assigned with a constant regression function within a given small window around the threshold (Cattaneo_Idrobo_Titiunik_2024), whereas we consider a smooth but nonconstant regression function within the window. For example, Armstrong.Kolesar2018 and Imbens.Wager2019 propose finite-sample minimax linear estimators under the smoothness of the regression function. However, these minimax estimators require knowledge of the conditional variance function, which is generally unavailable in practice. Although the variance can be estimated, we cannot guarantee the theoretical validity of the plug-in estimators with the estimated variance in finite samples. Furthermore, the construction of finite-sample valid confidence intervals based on these estimators additionally requires the normality of the regression errors.
@@
In this study, we propose finite-sample estimation and inference methods for RD designs with binary outcomes. For a binary dependent variable, all features of its conditional distribution, including its conditional variance, are known functions of its conditional mean function. We establish the finite-sample validity of our methods under a smoothness restriction on the conditional mean function by considering the implicit restrictions imposed on the entire conditional distribution. Therefore, our procedure is both feasible and theoretically valid without knowledge or estimation of the conditional variance or, more generally, any features of the conditional distribution, except for the smoothness of the conditional mean.
@@
Specifically, we consider a minimax optimal estimator among a class of linear shrinkage estimators for the regression function at a boundary point, under the assumption that the regression function satisfies Lipschitz continuity. The class of linear shrinkage estimators has the form $\sum_{i=1}^{n}w_{i}(Y_{i}-1/2)+1/2$ with $\sum_{i=1}^{n}w_{i}\leq 1$ and $w_{i}\geq 0$, where $Y_{1},...,Y_{n}$ are the observed outcomes on either side of the boundary. The shrinkage toward $1/2$ is motivated by the fact that the regression function is bounded and takes values in $[0,1]$, leading to a scope of efficiency gain by shrinkage. Given this class of linear shrinkage estimators, we derive a linear shrinkage estimator that minimizes the maximum mean squared error (MSE) under Lipschitz continuity with a known Lipschitz constant. In other words, we assume the researcher’s a priori knowledge of the bound of how much the function value can change if the running variable is changed by one unit. We emphasize that the Lipschitz constant is the only tuning parameter used. Furthermore, we show that the minimax estimator is a solution to a convex optimization problem that is computationally feasible. Thus, we provide a practical estimator that achieves finite-sample optimality in the binary-outcome setting.
@@
Our estimator is applicable to many practical RD designs. Binary outcomes are among the most common types of outcomes in empirical applications. For example, the following outcome variables are all binary: an indicator for winning the next election in the famous U.S. House election study by Lee2008; a corruption indicator in Brollo.Nannicini.Perotti.Tabellini2013; a mortality indicator in card_does_2009; and indicators for student’s enrollment and dropout in Melguizo2016 and Cattaneo2021multi. Furthermore, the first stage in fuzzy RD designs often involves treatment status as a binary dependent outcome. Moreover, the minimax optimality of our estimator for binary outcomes immediately extends to that for bounded outcomes because the variance of any linear estimator is maximized when the outcomes are Bernoulli, given the conditional mean function. Hence, our estimator can be applied not only to binary outcomes but also to bounded outcomes, which are also frequently used in the RD design.
@@
Our method also complements existing minimax estimators. We compare our estimator to a version of the existing minimax estimators (Armstrong.Kolesar2018; Imbens.Wager2019) and demonstrate that our method has better finite-sample performance than the existing approach, although their asymptotic behaviors are similar. Specifically, we consider a minimax linear estimator obtained under a misspecified model in which the conditional mean and variance are unrelated, the variance is known, and the regression function lies in a Lipschitz class with no bounds on the function values. This estimator is not directly feasible in our binary-outcome setting, in which the variance is unknown. As a feasible version of this estimator, we consider one obtained under the assumption of a constant variance of $1/4$, which is the maximum possible variance of a binary variable. For binary outcomes, we theoretically show that the efficiency gain from our estimator, relative to the alternative estimator above, tends to vanish as the sample size increases. Nevertheless, for small samples, we numerically demonstrate that the alternative method can result in a $5\%$ to $20\%$ increase in the worst-case root MSE owing to model misspecification. Hence, our method supplements the existing minimax estimators with better finite-sample performance and similar asymptotic behaviors in a binary-outcome setting.
@@
We also propose confidence intervals with correct coverage in finite samples uniformly over the Lipschitz class. We construct confidence intervals by inverting one- or two-sided uniformly valid tests that use a linear estimator as the test statistic. To construct a uniformly valid test, we propose a simulation-based approximation to the distribution of the test statistic by drawing samples from a multivariate Bernoulli distribution that satisfies the null restriction. We then numerically optimize the critical value so that the worst-case rejection probability is equal to or smaller than the significance level. A computational challenge with this approach is the calculation of the worst-case rejection probability, which involves the optimization of an $(n+1)$-dimensional parameter. We overcome this challenge by deriving a simple characterization of the worst-case rejection probability under Lipschitz continuity, which significantly reduces the computational burden. We also emphasize that our confidence intervals are valid in finite samples for binary outcomes. This contrasts with existing inference methods, which are based on either a large-sample approximation or a restrictive assumption of Gaussian errors with a known variance.
@@
The same inference approach does not apply to bounded outcomes because the simple characterization of the worst-case rejection probability relies on the fact that the outcome is binary. For bounded outcomes, we provide an alternative finite-sample inference procedure based on a uniform bound on the rejection probability obtained using Hoeffding’s inequality. The resulting confidence intervals have correct coverage in finite samples but can be conservative, similar to Hoeffding’s inequality-based confidence intervals in other contexts.
@@
We demonstrate the performance of our methods through simulations and an empirical application. In the simulations, our estimator achieves substantially smaller MSEs relative to the leading large-sample estimators when the sample size is small. Furthermore, our estimator behaves similarly to the large-sample estimators when the sample size is large; the differences in MSE decrease as the number of observations increases. Our proposed inference method also achieves guaranteed coverage rates with shorter confidence intervals when the sample size is small. Hence, our methods, while theoretically valid, are also useful in practice.
@@
We illustrate our methods by revisiting Brollo.Nannicini.Perotti.Tabellini2013, who estimate the impact of additional government revenue on corruption. They exploit the regional fiscal rule in Brazil, where federal transfers to municipal governments change exogenously at given population thresholds. This is a multi-cutoff RD design with a small sample size near each cutoff point. We demonstrate that our estimates are similar to the conventional estimates for large-sample pooling of multiple cutoffs. Nevertheless, our inference method provides much shorter confidence intervals than conventional methods when we focus on a small sample near each cutoff value. Consequently, our estimates provide more informative results than conventional methods.
@@
Both the simulation and application results indicate that the small-sample estimations are generally challenging, whereas our estimator has the potential to provide informative estimates. Hence, our estimator is a practical last resort for an empirical researcher facing a research question with a small effective sample size for an RD design.
@@
In addition to contributing to estimation in RD design, we contribute to the vast literature on minimax estimation. donoho1994 considers minimax affine estimation and inference on linear functionals in nonparametric regression models with Gaussian errors. Recently, this framework has been applied to the estimation and inference of treatment effects in various settings, including RD designs (Armstrong.Kolesar2018; Armstrong2021ATE; Gao2018; Imbens.Wager2019; Kwon.Kwon2020; Chaisemartin2021; rambachan2023parallel). We complement these studies by examining nonparametric regression models with Bernoulli dependent variables, which are not covered by their frameworks. To the best of our knowledge, no general minimax estimator under the squared error loss has been established for the problem of estimating linear functionals in this setting. [^2] DeRouen1974 derive a $\Gamma$-minimax estimator for a linear combination of the success probabilities of multiple independent binomial variables when the class of prior distributions consists of distributions with the same, known means. No solution is known, even for the estimation of the difference in the success probability between two independent binomial variables with unequal numbers of trials (Lehmann1998, Example 5.1.9). [^3] For the estimation of the success probability of a single binomial variable, a linear shrinkage (toward $1/2$) estimator is minimax among all estimators (Lehmann1998, Example 5.1.7). Marchand2000 consider this problem with a restricted parameter space. They show that, when the success probability is known to lie in a symmetric interval around $1/2$, a linear shrinkage estimator is minimax among all linear estimators. We contribute to this underexplored literature by developing a minimax estimator for a regression function at a point within the class of linear shrinkage estimators under the Lipschitz continuity of the regression function.
@@ 2 Our minimax estimator and its properties
RD designs exploit a discontinuous change in treatment status when a running variable exceeds a cutoff point. For example, Brollo.Nannicini.Perotti.Tabellini2013 exploit discontinuous increases in the amount of central government subsidies for a local government when its population equals or exceeds a threshold level. The target parameter of the RD design is the average treatment effect at the cutoff point, which is identified as the difference between the conditional expectation functions evaluated at the cutoff point. Hence, its estimation involves the nonparametric estimation of the conditional mean functions at their boundary points.
@@ 2.1 Setting
Suppose that we have a random sample $\{Y_{i},D_{i},R_{i}\}_{i=1}^{N}$, where $R_{i}\in\mathbb{R}^{d_{r}}$ is a $d_{r}(\geq 1)$-dimensional vector of running variables, $Y_{i}$ is a binary outcome, $D_{i}$ is a binary treatment assigned as $D_{i}=1\{R_{i}\in\mathcal{T}\}$, and $\mathcal{T}\subset\mathbb{R}^{d_{r}}$ is a known treated region. The leading case is that in which $R_{i}$ is univariate ($d_{r}=1$) and $\mathcal{T}=[c,\infty)$ for some known cutoff $c$; however, the following arguments also apply to a multidimensional case (i.e., $d_{r}>1$). Suppose
@@
for some unknown function $f:\{0,1\}\times\mathbb{R}^{d_{r}}\rightarrow[0,1]$. Let $R_{0}$ be a fixed boundary point in the treatment region $\mathcal{T}$. When $f(d,r)$ represents the conditional expectation function of the underlying potential outcome $Y_{i}(d)$ conditional on $R_{i}=r$ for each $d\in\{0,1\}$, $f(1,R_{0})-f(0,R_{0})$ is interpreted as the average treatment effect at the boundary point $R_{0}$ (hahnIdentificationEstimationTreatment2001). The data $\{Y_{i},D_{i},R_{i}\}_{i=1}^{N}$ can be divided into $\{Y_{i,+},R_{i,+}\}_{i=1}^{n_{+}}$ and $\{Y_{i,-},R_{i,-}\}_{i=1}^{n_{-}}$, where the former is the data from the treatment group and the latter is the data from the control group. We use the two samples separately to estimate $f(1,R_{0})$ and $f(0,R_{0})$, respectively.
Without loss of generality, we consider the estimation of $f(1,R_{0})$ throughout this section, except in Remark 2.6 at the end of this section, where we discuss the estimation of $f(1,R_{0})-f(0,R_{0})$. To simplify the notation, we use $\{Y_{i},R_{i}\}_{i=1}^{n}$ to denote $\{Y_{i,+},R_{i,+}\}_{i=1}^{n_{+}}$, so that $R_{i}\in\mathcal{T}$ for all i=1,ni=1,...n. Furthermore, we use $f(\cdot)$ to denote $f(1,\cdot)$. Additionally, our analysis conditions on the realization of $\{R_{i}\}_{i=1}^{n}$, and we treat $\{R_{i}\}_{i=1}^{n}$ as deterministic, so that $P(Y_{i}=1)=f(R_{i})$ for all $i=1,\ldots,n$. Let $p_{i}\equiv f(R_{i})$ for $i=0,1,\ldots,n$ and $\bm{p}\equiv(p_{0},p_{1},\ldots,p_{n})^{\prime}\in[0,1]^{n+1}$. Without loss of generality, we assume that $R_{0}=0$ and $\|R_{0}\|\leq\|R_{1}\|\leq\cdots\leq\|R_{n}\|$, where $\|\cdot\|$ is a norm on $\mathbb{R}^{d_{r}}$. The following theoretical result holds for any norm. We focus on the Euclidean norm in our numerical exercises, simulations, and empirical application.
For the parameter of interest $p_{0}=f(R_{0})=f(0)$, we consider the following linear shrinkage estimator:
@@
where $\mathcal{W}\equiv\left\{\bm{w}\in\mathbb{R}^{n}:\sum_{i=1}^{n}w_{i}\leq 1\ \text{and}\ w_{i}\geq 0\ \text{for all $i$}\right\}$. When $\sum_{i=1}^{n}w_{i}=1$, $\hat{p}_{0}(\bm{w})=\sum_{i=1}^{n}w_{i}Y_{i}$, and no shrinkage occurs. When $\sum_{i=1}^{n}w_{i}<1$, $\hat{p}_{0}(\bm{w})$ is an estimator that shrinks toward $1/2$.
We assume that $f$ belongs to a Lipschitz class:
@@
where $C$ denotes the Lipschitz constant and is known. This assumption implies that $\bm{p}\in[0,1]^{n+1}$ satisfies $|p_{i}-p_{j}|\leq C\|R_{i}-R_{j}\|$ for all $i$ and $j$. Conversely, if $|p_{i}-p_{j}|\leq C\|R_{i}-R_{j}\|$ for all $i$ and $j$, we can find a function $f\in\mathcal{F}_{\text{Lip}}(C)$ such that $f(R_{i})=p_{i}$ for all $i$ (Beliakov2006). Hence, the parameter space of $\bm{p}$ can be expressed as follows:
@@
To simplify the expression in (2.3), we redefine $p_{i}$ as $\theta_{i}\equiv p_{i}-1/2$ for $i=0,1,\ldots,n$ and let $\bm{\theta}\equiv(\theta_{0},\theta_{1},\ldots,\theta_{n})^{\prime}$; thus, the problem is
@@
Hence, we aim to obtain a weight vector that minimizes the maximum MSE by solving (2.4).
@@
The class of linear shrinkage estimators (2.1) eliminates linear estimators with negative weights. Hence, it excludes local polynomial estimators (except for local constant estimators), which are commonly employed in RD designs. Nevertheless, the linear minimax MSE estimator has nonnegative weights in related setups in which the outcome is nonbinary (e.g., Gaussian outcomes) and its regression function lies in the Lipschitz class with a known conditional variance: see Section 3 and Appendix D. Hence, we focus on linear shrinkage estimators with nonnegative weights.